Smith-Kline Company maintains inventory records at selling prices as well as at cost. For 2021, the records indicate the following data:

($ in 000s)
Cost Retail
Beginning inventory $ 105 $ 175
Purchases 833 1,097
Freight-in on purchases 40
Purchase returns 1 3
Net markups 5
Net markdowns 9
Net sales 926


Required:
Assuming the price level increased from 1.00 at January 1 to 1.60 at December 31, 2021, use the dollar-value LIFO retail method to approximate cost of ending inventory and cost of goods sold. (Do not round intermediate calculations. Round final answers to the nearest whole dollar. Enter your answers in thousands.)

Estimated Ending Inventory at retail

Estimated Ending inventory at cost

Estimated cost of goods sold
Answer #1

Cost $105 $832 $40 Retail $175 $1,094 (59) 2 Beginning inventory 3 Add: Net purchases 4 Freight-in 5 Net markups 6 Less: Net

Calculations are shown below: Cost 105 =833-1 40 Retail 175 =1097-3 -9 2 Beginning inventory 3 Add: Net purchases 4 Freight-i